How to solve tan225-cot81.cot69 whole divided by cot261+tan21

Grade:12th pass

1 Answers

manish
7 years ago
Tan225-cot81cot69/cot261+tan21=tan 45-tan9tan21/tan9+tan 21=1-tan9tan21/tan9+tan21=1/tan(9+21)=1/tan30=root 3Manish kumar panda